Charlie is half as old as Ramon, while Jun is 11 year older than Charlie. If half of Ramon's age 4 years from now will be 9 years less than the sum of half of Jun's age 4 years from now and charlie's present age. Find their ages
EQUATION:
1/2(x+4)=1/2(1/2x+15)+1/2x-9
Thanks
-Mark

Answer by richwmiller(17219)

c=r/2,
j=c+11,
1/2(r+4)=1/2*((j+4)+c)-9,
1/2(r+4)=1/2*((r/2+11+4)+r/2-9,
1/2(r+4)=1/2*(r/2+11+4)+r/2-9,
1/2(r+4)=1/2*(r/2+15)+r/2-9
r=14
c=7
j=18
which is your equation
1/2(x+4)=1/2(1/2x+15)+1/2x-9
x=14
check
1/2*(18)=1/2*22+7-9
9=11+7-9
9=9
ok